FEL Upgrade Project Weekly Brief
May 21-25, 2001





Highlights:
Excellent progress was made this week on diagnostic instrumentation and the dipole magnet design efforts for the
FEL upgrade.

The contract for the corrector dipole magnets has been placed.

We wish all our friends and colleagues a nice Memorial Day Holiday!

Management:
We continued preparations for the JTO sponsored workshop on High-Power FELs which will be held on June 5
and 6 in Newport News, VA, and the follow-up JTO FEL Working Group Meeting on June 7.

WBS 4 (Injector):
Conditioned the nitrogen implanted SST electrode to 12.5 Mv/m, the performance is not very good so far
emitting about 2 uA's at 125 kV.  Made corrections to cathode stalk drawings and continued work on our shield
door test setup.  The shield parts have been machined and are awaiting welding.

WBS 6 (RF):
Zone 1 - The brushes for unit 3 were replaced this week.  Testing of Zone 1 automatic ramp down voltage not
done, due to no resolution on how to lock up the tunnel.

Zone 2 - Operational.

Zone 3 - No progress.  Waiting for additional parts.

Zone 4 - The testing of the seven (7) klystrons heater voltage is complete.  The results are
not yet available.  In process of removing current 2.5 watt amps and replacing with 5 watt amps pending
resolution of ODH situation.

WBS 8 (Instrumentation):
Optical BPMs are making great progress. All testing has been done using the drive laser and turning mirrors on
picomotors.  The Auto zero function works very well; this is used once the optical transport has been setup the
detector is moved to zero rather than moving the optical transport to the same position for each lasing setup. The
transport lock chassis has been checked out, the last remaining thing is the Digital-to-Analog (D/A) converter.
This has real time control over six axis of transport mirrors (X & Y for Coll., can 2 & 6). The last detectors from
Molectron have not yet arrived.

The final wiring of the Lab #3 & #6 hutches was completed. Thanks to all who helped out including Jody Brock
for last minute welding.  Joe Gubeli had good comments regarding the failsafe operation of the different
combinations of turning mirrors and shutters.  These and the rest of the LSS will have the low & high level
certifications done over the next 10 days.

The 3" beam instrumentation devices keep coming in; 18 of the 36 shielded beam viewers are here and 32 of 44
3" beam position monitors are here.  The new "silicon" viewers are installed.  A new one was installed at
4F01while the one at 2F09 only had the screen changed from Al to Si; both were fitted with new cameras. The
air must still be connected to 4F01.  This activity included building and documenting a new camera/lamp power
supply box.

Drawings completed:  Ion Pump Power Supply Wiring Diagram, Ion Pump Power Supply Front PCB Schematic
and Assembly Drawings, Upstairs Dump Raster Board Schematic, General Purpose 64 Bit Digital I/O Card
Schematic and Assembly Drawings.

Drawings at EECAD:  Picomotor Relay Chassis Wiring Diagram, Beamviewer Camera/Light Board Power
Supply Box Wiring Diagram, Analog Differential Driver Board Schematic, Pyro-Electric Detector
Buffer/Amplifier, VME Timing Module Schematic, Fabrication and Assembly; Machine Protection
System-System Drawing (partial).

WBS 9 (Transport):
Dipoles
Injector Dipoles (DU/DV)
o  DULY Research, using small corrections to their magnetic model matched the core field and integrals of both
    the small and large Injector Dipoles at the same current, all within the tight 1 part in 10,000 of the required by
    Dave Douglasí specifications.
o  They continue work on the drawings of the DV to obtain a practical hardware design with its extra trim coil
     and sliders on the field clamp gap.
o  The drawings of the large injector dipole (GU) are being checked.
Arc Dipoles (GY, GX, GQ)
o  The magnetic analysis of the GY was successful in reducing the local field perturbation caused buy the large
    1.5 inch diameter holes for the diagnostic periscopes.  They extended the silicon steel around the hole.  They
    were also successful in widening the good field region to within specification by changing the inner edge of the
    silicon steel.  They will try a new series of runs with the model with reduced strength path length correctors.
    The smaller required slots should reduce the perturbation to the field so the horizontal parallel offset of the
    beam reduces to sub mm values. This is probably the final solution for this magnet.
o  We finished checking the drawings of the GY magnet and sent them to AES for back checking.
o  They are continuing design and magnetic modeling the Arc Bend Dipole (GX).  They removed the nubbins on
    the edges of the poles and obtained the 1 part in 10,000 bulk field and integral uniformity required of the
    specifications. However they have to work on raising the bulk field.  They also activated just the two of the four
    coils that face the gap in order to investigate the half-field switch-to-UV configuration.  With this they were also
    successful in obtaining the appropriate field uniformities.  This success allowed continued design of the current
    leads for the coils to enable this form of powering.  At the same time, the assembly drawing and detail
    drawings were firmed up.
Quadrupoles
QG (3 inch quad)
o  The coil-manufacturing package remains in Procurement cycle, having been extended by a week at the request
    of one vendor.
o  The core procurement package was sent out for bid.
o  We tried two added shim configurations to the prototype quadrupole in an attempt to understand what we
    could do to further reduce the icosapole, which has been rising as we increased the chamfer size.
QX (Trim Quad)
o  We reduced the length of the magnetic model by one inch and obtained a better overall quadrupole gradient
    integral.  The reduction was indicated to back the coils away from the field clamps of the GQ dipole.
Corrector Dipoles (DB & DJ)
o  We placed the contract with Milhous Co. of Lynchburg VA after they clarified their manufacturing plan.
Sextupole
o  Robin Wines continued converting her 2D design to 3D design in TOSCA.
General
o  We are working with the company that gave us the Dipole Steel bid to clarify details of their proposal.
o  We completed design layout of the girders between cryomodule such that we can hole a working session next
    week that fixes the design of this position and the return leg and allows detail drawings to start.
o  The Task definition to be used by an engineering service contractor for design of the arc chambers continues in
    checking.  Robby Hicks continues putting together the attachments and graphics.
o  The heavy-duty alignment cartridges from CERN (European Organization for Nuclear Research) for use under
    the large 180 degree dipoles are being shipped.

WBS 11 (Optics):
The design on the optical cavity assembly internals continues.  We worked on better defining the actual
components for the linear actuation.  We held discussions with the makers of lvdts and motorized actuators.  We
should receive test articles of both components next week, and will do outgassing studies.

The rest of the beampath shields for the mirror test stand arrived, and we will complete installation after we finish
some items required to begin Operations.  We received the modified IR Demo mirror mount, so we can better
quantify a number of parameters associated with our edge-cooled scheme.  These will be needed for the
outcoupler mirror on the Upgrade.

The Drive Laser was completely realigned with the new Nd:YLF laser rod.  Output as well-above
specifications.  The optical transport system (OTS) for the drive laser was realigned a far as the beam aperture.
We will complete some work on the RF PLL today for operations.

I attended a Joint Technical Office Technical Area Working Group , held in Albuquerque the 23rd and 24th.
While there, I gave presentations on laser/materials interactions work done here, and on the User Facility itself.
Attendees were quite interested in the unique combination of high average power and ultrashort pulsewidth.  The
workshop also gave me an opportunity to discuss support by staff from NAVSEA (Dahlgren) and NRL for the
series of experiments we'll do under the JTO contract.

Operations/Commissioning:
The following work was done this week to prepare for operations for the summer run:  Installed insertable
mirrors in U.L. 3 to divert beam from the metals processing hutch to the nanotubes experimental setup.  Installed
manual beam stops in U.L. 3.  Aligned FEL OTS.  Installed insertable mirror in U.L. 2 beam dump assembly.
Installed OD 1 filters in OBPMs (FL02 and FL08).  Aligned optics on OBPM test stand in the Drive Laser
Enclosure.

Work in the FEL building was temporarily suspended for one day when engineers performing a CEBAF helium
spill test discovered that the oxygen monitors do not respond correctly in a helium rich atmosphere.  These
sensors were our first line of warning for oxygen deficiency hazards and serve the same function at a number of
laboratories around the world.  It was a surprise to both the manufacturer and the scientific community that these
did not work as advertised. The FEL was able to re-establish normal operations through a series of mitigating
steps including locking open roof vents and removing doors to eliminate the hazards in most work areas.  Work
near the ceiling and in racks connected to the tunnel remains ODH-2 classification (requiring SCBA gear and a
two man rule) until such time as replacement monitors that work are installed.
 
 



Privacy and Security Notice